# Cosmic Ray Magnetohydrodynamics: A New Two-Moment Framework with Numerical Implementation

### v1

@@ 4 Numerical Tests of the CR subsystem
Figure 2: Numerical results of the 1D streaming test using the same setup as in Figure 1, except for elevated scattering coefficients $\sigma_{0,\rm st}$ to illustrate the performance of our stability criterion. Time steps $\Delta t$ progressively decrease from left to right, as indicated by the titles. We see numerical instability and deviation from analytical solutions persist until the stability condition $\Delta t<\sqrt{3}\Delta x/(1+4\sigma_{0,\rm st})V_{m}$ is satisfied in the rightmost column.
4.2 1D CR Streaming (non-linear Landau damping)
CR transport is also sensitive to the wave damping process, which balances wave growth, modulates CR scattering, and ultimately determines the level of coupling between CRs and the MHD gas. In our formulation, the effects of various damping mechanisms are essentially encapsulated into scattering coefficients. In this subsection, we test scattering coefficients of the form
$\sigma_{\pm}=\sigma_{0,\rm nll}\sqrt{\left|\nabla P_{\rm cr}\right|}\Theta(\mp\nabla P_{\rm cr}\cdot\vec{b})$ (70)
which correspond to the case where MHD waves are generated by CR streaming instability but damped by nonlinear Landau mechanism (Armillotta et al., 2022). Here $\sigma_{0,\rm nll}$ is a constant accounting for other environmental parameters, but with a dimension of ${\rm length}^{-1/2}\cdot{\rm velocity}^{-1}\cdot{\rm pressure}^{-1/2}$. The steady-state $F_{\rm cr}$ in this case (taking the gas velocity to be zero) is solved to be:
$F_{\rm cr}=-\left(\frac{\sqrt{|\nabla P_{\rm cr}|}}{\sigma_{0,\rm nll}}+\frac{4}{3}v_{A}\mathcal{E}_{\rm cr}\right){\rm sgn}(\nabla P_{\rm cr}\cdot\vec{b})$ (71)
We set up a 1D test similar to that in Section 4.1, with the initial $\mathcal{E}_{\rm cr}$ profile being:
$\mathcal{E}_{\rm cr}(0,x)=\frac{1}{\mathcal{E}_{0}^{-1}+3\sigma_{0,\rm nll}^{2}v_{0}^{2}|x|}$ (72)
where $\mathcal{E}_{0}$ and $v_{0}$ are constant parameters and $x\in(-L_{\rm box}/2,L_{\rm box}/2)$. One can find initially $\sqrt{|\nabla P_{\rm cr}|}=\sigma_{0,\rm nll}v_{0}\mathcal{E}_{\rm cr}$, such that $F_{\rm cr}=-\left(v_{0}+4v_{A}/3\right)\mathcal{E}_{\rm cr}{\rm sgn}(\nabla P_{\rm cr}\cdot\vec{b})$. Assuming $v_{A}$ and $v_{0}$ to be constant, this form of $F_{\rm cr}$ makes traveling waves the solutions to the $\mathcal{E}_{\rm cr}$ equation, and the fronts propagate at the speeds of $\pm(v_{0}+4v_{A}/3)$.
@@
$\frac{x_{p}}{(3\sigma_{0,\rm nll}^{2}v_{0}^{2}\mathcal{E}_{0})^{-1}+x_{p}-(v_{0}+4v_{A}/3)t}$ (75)
$={\rm log}\left[1+\frac{x_{p}-(v_{0}+4v_{A}/3)t}{(3\sigma_{0,\rm nll}^{2}v_{0}^{2}\mathcal{E}_{0})^{-1}}\right]$
In our test problem, we set $\sigma_{0,\rm nll}=2$, $v_{0}=0.5$, $\mathcal{E}_{0}=0.1$ and $L_{\rm box}=4$. The MHD gas properties are identical to those in Section 4.1, so that $v_{A}=1$. We compare the reference solution Equation 73 with numerical results in Figure 3, and see excellent agreement. Therefore, our code can incorporate various types of scattering coefficients which essentially correspond to different microphysical processes.
We comment that our stability criterion, although not designed for this test, in practice still works fine with a reduction factor of $\sim 2.35$. Directly equating Equation 68 and Equation 70, the effective $\sigma_{0,\rm st}^{*}$ in this test is about 0.67. Substituting it directly back to the stability criterion Equation 66 gives the stable time step $\Delta t=0.47\Delta x/V_{m}$, while we roughly achieve stability with $\Delta t=0.2\Delta x/V_{m}$ used in the test above. We used a spatial resolution $\Delta x=1/64$ here. However, we caution that, details of numerical requirements of different forms of scattering coefficients (particularly on the time step) await future explorations.
Figure 3: Numerical results of 1D streaming tests from Section 4.2. This is similar to Figure 1 but the damping mechanism follows the form expected for non-linear Landau damping, with the scattering coefficients defined in Equation 70. Dashed lines denote our theoretical predictions, which align well with the simulation data at the propagating fronts. Deviations occur in the central plateau, where $\nabla P_{\rm cr}\rightarrow 0$ violates our theoretical assumption of a non-negligible $\nabla P_{\rm cr}$. As a result $\mathcal{E}_{\rm cr}$ and $F_{\rm cr}$ smoothly transition to the propagating fronts.
4.3 1D CR Diffusion
@@
To test this transport regime, we set all the four scattering coefficients constant and equal: $\sigma_{+}^{L}=\sigma_{+}^{R}=\sigma_{-}^{L}=\sigma_{-}^{R}\equiv\sigma_{0,\rm diff}/4$ with $\sigma_{0,\rm diff}$ denoting their summation. By solving for the steady $F_{\rm cr}$ and plugging it back into the CR energy Equation 67, we obtain
@@
where gas velocity $u$ is assumed uniformly constant. Combining the first two terms as a comoving derivative $D\mathcal{E}_{\rm cr}/Dt$, we see it becomes a standard diffusion equation in the frame moving [^7] The fact that $\mathcal{E}_{\rm cr}$ is advected at a different velocity from the background gas breaks Galilean invariance. The reason is that by neglecting the source term in equation Equation 67 we ignored the work done by $P_{\rm cr}$ on the moving gas, and thus reach to a somewhat unphysical situation. After recovering the $\mathcal{E}_{\rm cr}$ source term, one can see in reality the $\mathcal{E}_{\rm cr}$ profile should flow at the same velocity as the background gas (provided the $v_{A}^{2}$ heating terms corresponding to the Fermi acceleration are omitted). with the velocity $4u/3$. We initialize the system with a Gaussian profile $\mathcal{E}_{\rm cr}(0,x)={\rm exp}(-40x^{2})$, and the theoretical solution to the diffusion equation can be obtained via Green’s function method. When the background gas is static, Jiang & Oh (2018) wrote down the analytical result:
@@
Figure 4: Numerical results of 1D diffusion tests described in Section 4.3. Top panels: Evolution of $\mathcal{E}_{\rm cr}$ profiles in a static MHD background. Bottom panels: Evolution of $\mathcal{E}_{\rm cr}$ profiles in a uniformly moving MHD background (MHD velocity u=1u=-1).
@@ 4.4 Anisotropic CR Flux
In this subsection, we test the performance of the Lorentz term with a commonly employed problem setup (e.g., Parrish & Stone, 2005; Sharma & Hammett, 2011; Pakmor et al., 2016; Jiang & Oh, 2018), where the magnetic field is circular in a 2D Cartesian $(x,y)\in(-1,1)\times(-1,1)$ domain:
@@
where $r\equiv\sqrt{x^{2}+y^{2}}$ and $\phi={\rm atan2}(y,x)$ is the angle between the positive $x$-axis and the ray from the origin to the point $(x,y)$. In this test we arrange both $\sigma_{+}$ and $\sigma_{-}$ constant regardless of the sign of $\nabla P_{\rm cr}$, so that the steady state flux equation becomes $\nabla P_{\rm cr}=-\left(\sigma_{+}+\sigma_{-}\right)F_{\rm cr}$ and $\mathcal{E}_{\rm cr}$ is subject to a traditional diffusion equation. The initial ${F}_{\rm cr}$ is set zero. Under this circumstance, Pakmor et al. (2016) has derived the analytical solution to compare with
@@
Our numerical results are compared with the analytical solution (bottom right) in Figure 5. All simulations exhibit the expected anisotropic transport. While lower-resolution runs show increased numerical diffusion, the $1024\times 1024$ run agrees closely with the analytical solution. The spatial convergence here is consistent with the corresponding test in Jiang & Oh (2018).
Figure 5: 2D tests for the anisotropic transport of $\mathcal{E}_{\rm cr}$ along magnetic field lines on Cartesian grids, as described in Section $4.$ An initial patch with high CR energy density streams along circular background magnetic fields. The bottom right panel shows the analytical solution while other three panels demonstrate numerical results with different spatial resolutions at the same time.
4.5 CR Streaming in Spherical Coordinates
@@
We use a CR streaming problem similar to the one described in Section 4.1, except we now run it on a spherical grid with $512\times 32\times 8$ cells in the $r$, $\theta$ and $\phi$ directions, respectively. This time, we initialize the $\mathcal{E}_{\rm cr}$ profile by setting $\mathcal{E}_{\rm cr,0}=2$ in the central core $r<r_{\rm core}\ (r_{\rm core}=1.0)$, and 0 in the rest of the simulation domain. According to Equation 69, the initial discontinuous front should propagate at the speed $v_{\rm front}=(1+4\sigma_{0,\rm st})v_{A}/3\sigma_{0,\rm st}$, where we set $\sigma_{0,\rm st}=1$. For the purpose of this test we impose a monopole magnetic field so that the CR streaming is only along the radial direction, and we arrange the gas density accordingly to keep $v_{A}=1$ also a constant. Although a monopole magnetic field is unphysical, here we turn off the MHD evolution and the setup for gas only serves as a static platform.
@@
Figure 6: Numerical results of the CR streaming problem in a 3D spherical coordinate system, as described in Section 4.5, with CR transport along the radial direction. Top panels: $\mathcal{E}_{\rm cr}$ profiles shown in $r-\theta$ slices at different times. The boundary separating the central high-$\mathcal{E}_{\rm cr}$ region and the outer zero-$\mathcal{E}_{\rm cr}$ region should expand at a constant speed $v_{\rm front}$, with $\mathcal{E}_{\rm cr}$ being constant in the central region, whose value can be calculated through CR energy conservation. Bottom panel: radial profiles of $\mathcal{E}_{\rm cr}$ compared with the theoretical expectation. We see they agree well and our code can properly handle the sharp discontinuity.
4.6 CRPAI in Cylindrical Coordinates
@@
We initialize $\mathcal{E}_{\rm cr}$ with a central peak, prompting CRs to stream outward along field lines. Consequently, as CRs traverse regions with varying magnetic field strengths, adiabatic invariance induces CR pressure anisotropy, activating the CRPAI. Following equations Equation 44 and Equation 45, in this scenario we prescribe the scattering coefficients and $\Delta P_{\rm cr}$ as:
$\Delta P_{\rm cr}=\frac{1}{\sigma_{0,\rm aniso}}\frac{v_{A}}{c}P_{\rm cr}$ (82)
$\sigma_{+}^{R}=\sigma_{-}^{L}=\frac{\sigma_{0,\rm aniso}}{2}\frac{1}{v_{A}c}\left|\frac{\dot{B}}{B}\right|$
$\sigma_{-}^{L}=\sigma_{-}^{R}=0$
where the dimensionless constant $\sigma_{0,\rm aniso}$ encapsulates the environmental parameters. In this test, we turn on the CR energy source term; otherwise the scattering coefficients in front of $\Delta P_{\rm cr}$ in the $\vec{F}_{\rm cr}$ source would cancel, allowing $\Delta P_{\rm cr}$ to influence evolution only through the energy source term.
@@
$=\left(\frac{4\sigma_{0,\rm aniso}-1}{3}\right)\frac{1}{B}\left|\frac{dB}{dR}\right|\mathcal{E}_{\rm cr}v_{A}$
@@
$=-\frac{\sigma_{0,\rm aniso}}{v_{A}}\frac{1}{B}\left|\frac{dB}{dR}\right|F_{\rm cr}.$
Generally speaking, the reference solution of $\mathcal{E}_{\rm cr}$ is a series of Bessel functions, and we leave the derivation in Appendix C. For simplification, we further set $\sigma_{0,\rm aniso}=0.25$ so that the CR energy source term vanishes. Physically it means we artificially balance Fermi acceleration with energy loss through CR heating. In this case, one reference solution of $\mathcal{E}_{\rm cr}$ is $\mathcal{E}_{\rm cr}(t,R)=J_{0}(\sqrt{3}R/2)e^{-t}$, where $J_{0}$ is the zeroth order Bessel function of the first kind. We use this theoretical prediction as initial condition, and prescribe it in the ghost zones, expecting an overall exponential decay $e^{-t}$ in the simulations.
@@
Figure 7: Evolution of $\mathcal{E}_{\rm cr}$ in a 2D cylindrical coordinate, where the CR scattering is subject to CRPAI, as described in Section 4.5. Top panels: 2D profiles of $\mathcal{E}_{\rm cr}$ at different times (title). We see the system is well symmetric in the angular direction. Bottom panel: radial profiles of $\mathcal{E}_{\rm cr}$ at the same moments as in the top panels. Simulations well match the expected exponential decay.
5 Numerical Tests of the full CR-MHD system
We now turn on the CR energy and momentum feedback to the MHD gas, validating our code for the coupled CR-MHD system. In doing so, we conduct the first linear analysis of the full two-moment CR-MHD wave system. The problem setup and full derivation/analysis are provided in Appendix A. Early investigations of CR-MHD waves by Begelman & Zweibel (1994) have revealed the presence of CR acoustic instability, which was revisited numerically by Tsung et al. (2022) recently. Our analysis focuses on the waves in a uniform fluid and CR background, but assumes the presence of a weak background $\nabla P_{\rm cr}$ to set the scattering rate, which is taken to be constant. Major highlights include:

### v2

@@ IV Numerical Tests of the CR subsystem
Figure 2: Numerical results of the 1D streaming test using the same setup as in Figure 1, except for elevated scattering coefficients $\widetilde{\sigma}_{0,\rm st}$ to illustrate the performance of our stability criterion. Time steps $\Delta t$ progressively decrease from left to right, as indicated by the titles. We see numerical instability and deviation from analytical solutions persist until the stability condition $\Delta t<\sqrt{3}\Delta x/(1+4\widetilde{\sigma}_{0,\rm st})V_{m}$ is satisfied in the rightmost column.
IV.2 1D CR Streaming (non-linear Landau damping)
CR transport is also sensitive to the wave damping process, which balances wave growth, modulates CR scattering, and ultimately determines the degree of coupling between CRs and the MHD gas (Armillotta et al., 2022; Thomas et al., 2025b). In our formulation, the influence of various damping mechanisms is essentially encapsulated into scattering coefficients. When CR scattering is mediated by MHD waves driven by CRSI but damped primarily through nonlinear Landau damping, the scattering coefficients take the form (e.g., Armillotta et al., 2022; Hix et al., 2025):
$\sigma_{\pm}^{L,R}=\frac{\widetilde{\chi}_{0,\rm NLL}}{2}P_{\rm therm}^{-1/4}\left(\frac{e\left|\nabla P_{\rm cr}\right|}{mc^{4}}\right)^{1/2}\Theta(\mp\nabla P_{\rm cr}\cdot\vec{b})$ (70)
where $e/m$ is the charge-to-mass ratio of a CR particle and $\widetilde{\chi}_{0,\rm NLL}$ is a dimensionless constant prefactor absorbing microphysical uncertainties or calibration.
In this subsection, we test the nonlinear Landau damping mechanism by performing a 1D test similar to that in Section IV.1, again holding the MHD variables fixed. Since $P_{\rm therm}=1$ is a constant, we group the associated environmental factors into a new dimensional constant $\chi_{0,\rm NLL}\equiv\widetilde{\chi}_{0,\rm NLL}P_{\rm therm}^{-1/4}(e/m)^{1/2}c^{-2}$ for numerical convenience. The resulting scattering coefficients are then
$\sigma_{\pm}=\chi_{0,\rm NLL}\sqrt{\left|\nabla P_{\rm cr}\right|}\Theta(\mp\nabla P_{\rm cr}\cdot\vec{b}).$ (71)
The steady-state $F_{\rm cr}$ in this case (taking the gas velocity to be zero) is solved to be:
$F_{\rm cr}=-\left(\frac{\sqrt{|\nabla P_{\rm cr}|}}{\chi_{0,\rm NLL}}+\frac{4}{3}v_{A}\mathcal{E}_{\rm cr}\right){\rm sgn}(\nabla P_{\rm cr}\cdot\vec{b})$ (72)
We initialize the $\mathcal{E}_{\rm cr}$ profile as:
$\mathcal{E}_{\rm cr}(0,x)=\frac{1}{\mathcal{E}_{0}^{-1}+3\chi_{0,\rm NLL}^{2}v_{0}^{2}|x|}$ (73)
where $\mathcal{E}_{0}$ and $v_{0}$ are constant parameters and $x\in(-L_{\rm box}/2,L_{\rm box}/2)$. One can find initially $\sqrt{|\nabla P_{\rm cr}|}=\chi_{0,\rm NLL}v_{0}\mathcal{E}_{\rm cr}$, such that $F_{\rm cr}=-\left(v_{0}+4v_{A}/3\right)\mathcal{E}_{\rm cr}{\rm sgn}(\nabla P_{\rm cr}\cdot\vec{b})$. Assuming $v_{A}$ and $v_{0}$ to be constant, this form of $F_{\rm cr}$ makes traveling waves the solutions to the $\mathcal{E}_{\rm cr}$ equation, and the fronts propagate at the speeds of $\pm(v_{0}+4v_{A}/3)$.
@@
$\frac{x_{p}}{(3\chi_{0,\rm NLL}^{2}v_{0}^{2}\mathcal{E}_{0})^{-1}+x_{p}-(v_{0}+4v_{A}/3)t}$ (76)
$={\rm log}\left[1+\frac{x_{p}-(v_{0}+4v_{A}/3)t}{(3\chi_{0,\rm NLL}^{2}v_{0}^{2}\mathcal{E}_{0})^{-1}}\right]$
In our test problem, we set $\chi_{0,\rm NLL}=2$, $v_{0}=0.5$, $\mathcal{E}_{0}=0.1$ and $L_{\rm box}=4$. The MHD gas properties are identical to those in Section IV.1, so that $v_{A}=1$. We compare the reference solution Equation 74 with numerical results in Figure 3, and see excellent agreement. Therefore, our code can incorporate various types of scattering coefficients which essentially correspond to different microphysical processes.
We comment that our stability criterion, although not designed for this test, in practice still works fine with a reduction factor of $\sim 2.35$. Directly equating Equation 68 and Equation 70, the effective $\widetilde{\sigma}_{0,\rm st}^{\rm NLL}$ in this test is about 0.67. Substituting it directly back to the stability criterion Equation 66 gives the stable time step $\Delta t=0.47\Delta x/V_{m}$, while we roughly achieve stability with $\Delta t=0.2\Delta x/V_{m}$ used in the test above. We used a spatial resolution $\Delta x=1/64$ here. However, we caution that, details of numerical requirements of different forms of scattering coefficients (particularly on the time step) await future explorations.
Figure 3: Numerical results of 1D streaming tests from Section IV.2. This is similar to Figure 1 but the damping mechanism follows the form expected for non-linear Landau damping, with the scattering coefficients defined in Equation 70. Dashed lines denote our theoretical predictions, which align well with the simulation data at the propagating fronts. Deviations occur in the central plateau, where $\nabla P_{\rm cr}\rightarrow 0$ violates our theoretical assumption of a non-negligible $\nabla P_{\rm cr}$. As a result $\mathcal{E}_{\rm cr}$ and $F_{\rm cr}$ smoothly transition to the propagating fronts.
IV.3 1D CR Diffusion
@@
To test this transport regime, we set all the four scattering coefficients equal and constant: $\sigma_{+}^{L}=\sigma_{+}^{R}=\sigma_{-}^{L}=\sigma_{-}^{R}\equiv\sigma_{0,\rm diff}/4$ with $\sigma_{0,\rm diff}$ denoting their summation. By solving for the steady $F_{\rm cr}$ and plugging it back into the CR energy density Equation 67, we obtain
@@
where gas velocity $u$ is assumed uniformly constant. Combining the first two terms as a comoving derivative $D\mathcal{E}_{\rm cr}/Dt$, we see it becomes a standard diffusion equation in the frame moving [^6] The fact that $\mathcal{E}_{\rm cr}$ is advected at a different velocity from the background gas breaks Galilean invariance. The reason is that by neglecting the source term in Equation 67 we ignored the work done by $P_{\rm cr}$ on the moving gas, and thus reach to a somewhat unphysical situation. After recovering the $\mathcal{E}_{\rm cr}$ source term, one can see in reality the $\mathcal{E}_{\rm cr}$ profile should flow at the same velocity as the background gas (provided the $v_{A}^{2}$ heating terms corresponding to the Fermi acceleration are omitted). with the velocity $4u/3$. We initialize the system with a Gaussian profile $\mathcal{E}_{\rm cr}(0,x)={\rm exp}(-40x^{2})$, and the theoretical solution to the diffusion equation can be obtained via Green’s function method. When the background gas is static, Jiang and Oh (2018) wrote down the analytical result:
@@
Figure 4: Numerical results of 1D diffusion tests described in Section IV.3. Top panels: Evolution of $\mathcal{E}_{\rm cr}$ profiles in a static MHD background. Bottom panels: Evolution of $\mathcal{E}_{\rm cr}$ profiles in a uniformly moving MHD background (MHD velocity u=1u=-1).
@@ IV.4 Anisotropic CR Flux
In this subsection, we test the performance of the Lorentz term with a commonly employed problem setup (e.g., Parrish and Stone, 2005; Sharma and Hammett, 2011; Pakmor et al., 2016; Jiang and Oh, 2018), where the magnetic field is circular in a 2D Cartesian $(x,y)\in(-1,1)\times(-1,1)$ domain:
@@
where $r\equiv\sqrt{x^{2}+y^{2}}$ and $\phi={\rm atan2}(y,x)$ is the angle between the positive $x$-axis and the ray from the origin to the point $(x,y)$. In this test we arrange both $\sigma_{+}$ and $\sigma_{-}$ constant regardless of the sign of $\nabla P_{\rm cr}$, so that the steady state flux equation becomes $\nabla P_{\rm cr}=-\left(\sigma_{+}+\sigma_{-}\right)F_{\rm cr}$ and $\mathcal{E}_{\rm cr}$ is subject to a traditional diffusion equation. The initial ${F}_{\rm cr}$ is set to zero. Under this circumstance, Pakmor et al. (2016) has derived the analytical solution to compare with
@@
Our numerical results are compared with the analytical solution (bottom right) in Figure 5. All simulations exhibit the expected anisotropic transport. While lower-resolution runs show increased numerical diffusion, the $1024\times 1024$ run agrees closely with the analytical solution. The spatial convergence here is consistent with the corresponding test in Jiang and Oh (2018).
Figure 5: 2D tests for the anisotropic transport of $\mathcal{E}_{\rm cr}$ along magnetic field lines on Cartesian grids, as described in Section IV.4. An initial patch with high CR energy density streams along circular background magnetic fields. The bottom right panel shows the analytical solution while other three panels demonstrate numerical results with different spatial resolutions at the same time.
IV.5 CR Streaming in Spherical Coordinates
@@
We use a CR streaming problem similar to the one described in Section IV.1, except we now run it on a spherical grid with $512\times 32\times 8$ cells in the $r$, $\theta$ and $\phi$ directions, respectively. This time, we initialize the $\mathcal{E}_{\rm cr}$ profile by setting $\mathcal{E}_{\rm cr,0}=2$ in the central core $r<r_{\rm core}\ (r_{\rm core}=1.0)$, and 0 in the rest of the simulation domain. According to Equation 69, the initial discontinuous front should propagate at the speed $v_{\rm front}=(1+4\widetilde{\sigma}_{0,\rm st})v_{A}/3\widetilde{\sigma}_{0,\rm st}$, where we set $\widetilde{\sigma}_{0,\rm st}=1$. For the purpose of this test we impose a monopole magnetic field so that the CR streaming is only along the radial direction, and we arrange the gas density accordingly to keep $v_{A}=1$ also a constant. Although a monopole magnetic field is unphysical, here we turn off the MHD evolution and the setup for gas only serves as a static platform.
@@
Figure 6: Numerical results of the CR streaming problem in a 3D spherical coordinate system, as described in Section IV.5, with CR transport along the radial direction. Top panels: $\mathcal{E}_{\rm cr}$ profiles shown in $r-\theta$ slices at different times. The boundary separating the central high-$\mathcal{E}_{\rm cr}$ region and the outer zero-$\mathcal{E}_{\rm cr}$ region should expand at a constant speed $v_{\rm front}$, with $\mathcal{E}_{\rm cr}$ being constant in the central region, whose value can be calculated through CR energy conservation. Bottom panel: radial profiles of $\mathcal{E}_{\rm cr}$ compared with the theoretical expectation. We see they agree well and our code can properly handle the sharp discontinuity.
IV.6 CRPAI in Cylindrical Coordinates
@@
We initialize $\mathcal{E}_{\rm cr}$ with a central peak, prompting CRs to stream outward along field lines. Consequently, as CRs traverse regions with varying magnetic field strengths, adiabatic invariance induces CR pressure anisotropy, activating the CRPAI. Following Equation 44 and Equation 45, in this scenario we prescribe the scattering coefficients and $\Delta P_{\rm cr}$ as:
$\Delta P_{\rm cr}=\frac{1}{\widetilde{\sigma}_{0,\rm aniso}}\frac{v_{A}}{c}P_{\rm cr}$ (83)
$\sigma_{+}^{R}=\sigma_{-}^{L}=\frac{\widetilde{\sigma}_{0,\rm aniso}}{2}\frac{1}{v_{A}c}\left|\frac{\dot{B}}{B}\right|$
$\sigma_{+}^{L}=\sigma_{-}^{R}=0$
where the dimensionless constant $\widetilde{\sigma}_{0,\rm aniso}$ encapsulates the environmental parameters. In this test, we turn on the $\mathcal{E}_{\rm cr}$ source term; otherwise the scattering coefficients in front of $\Delta P_{\rm cr}$ in the $\vec{F}_{\rm cr}$ source would cancel, allowing $\Delta P_{\rm cr}$ to influence evolution only through the CR energy source term.
@@
$=\left(\frac{4\widetilde{\sigma}_{0,\rm aniso}-1}{3}\right)\frac{1}{B}\left|\frac{dB}{dR}\right|\mathcal{E}_{\rm cr}v_{A}$
@@
$=-\frac{\widetilde{\sigma}_{0,\rm aniso}}{v_{A}}\frac{1}{B}\left|\frac{dB}{dR}\right|F_{\rm cr}.$
Generally speaking, the reference solution of $\mathcal{E}_{\rm cr}$ is a series of Bessel functions, and we leave the derivation in Appendix C. For simplification, we further set $\widetilde{\sigma}_{0,\rm aniso}=0.25$ so that the CR energy source term vanishes. Physically it means we artificially balance Fermi acceleration with energy loss through CR heating. In this case, one reference solution of $\mathcal{E}_{\rm cr}$ is $\mathcal{E}_{\rm cr}(t,R)=J_{0}(\sqrt{3}R/2)e^{-t}$, where $J_{0}$ is the zeroth order Bessel function of the first kind. We use this theoretical prediction as initial condition, and prescribe it in the ghost zones, expecting an overall exponential decay $e^{-t}$ in the simulations.
@@
Figure 7: Evolution of $\mathcal{E}_{\rm cr}$ in a 2D cylindrical coordinate, where the CR scattering is subject to CRPAI, as described in Section IV.5. Top panels: 2D profiles of $\mathcal{E}_{\rm cr}$ at different times (title). We see the system is well symmetric in the angular direction. Bottom panel: radial profiles of $\mathcal{E}_{\rm cr}$ at the same moments as in the top panels. Simulations well match the expected exponential decay.
V Numerical Tests of the full CR-MHD system
We now turn on the CR energy and momentum feedback to the MHD gas, validating our code for the coupled CR-MHD system. In doing so, we conduct the first linear analysis of the full two-moment CR-MHD wave system. The problem setup and full derivation/analysis are provided in Appendix A. Early investigations of CR-MHD waves by Begelman and Zweibel (1994) have revealed the presence of CR acoustic instability, which was revisited numerically by Tsung et al. (2022) recently. Our analysis focuses on the waves in a uniform fluid and CR background, but assumes the presence of a weak background $\nabla P_{\rm cr}$ to set the scattering rate, which is taken to be constant. Major highlights include: